OpenStudy (anonymous):

there is HC (hydrocarbon) on complete combustion gives 2.64 g CO2 and 0.54 g of H2O find out molecule of HC if the mass of 10 litre gas at ntp = 34.8 g

OpenStudy (abmon98):

Number of Moles=Mass(g)/Molar Mass 2.64/44.0=0.06 moles of CO2 0.54/18=0.03 moles of H2O Ratio of Moles of CO2 to H2O is 2:1 C2H2+5/2O2-->2CO2+H2O At STP 1 mole occupies 22.4 Litre V=22.4*Number of moles 10=22.4*34.8/Molar Mass Molar Mass:77.952 77.952/(12*2)+2=2.99 2.99*C2H2=C6H6